# The Remarkable Robustness of LLMs: Stages of Inference?

### v1

@@ 5.3 Stage 3: Prediction Ensembling
After about the halfway point, the model must begin converting semantic features into concrete predictions for the next token. Given the robustness observed in Figure 3, we hypothesize that this is accomplished with a kind of ensembling. Ensembling in neural networks with residual connections is akin to having many subnetworks perform a "vote" for the output [65]. Therefore, interfering with a single member of an ensemble is unlikely to have a destructive effect.
Prediction Neurons
@@
Previous work suggests that networks contain ensembles of “prediction" neurons, which act as probability promoters [66, 24, 32] and work in tandem with suppression neurons (Section 5.4). Following [32], we find prediction and suppression neurons by analyzing the output weights with the unembedding matrix . Prediction neurons exhibit a logit effect distribution with high kurtosis and positive skew, while suppression neurons show high kurtosis but negative skew. Here, is the output MLP weight for a given layer. Across 11 models, prediction neurons emerge around the midpoint, increasing in density towards the latter layers 6, before being outstripped by suppression neurons. To confirm their action, we study how much the model’s abstract representation changes as a function of prediction neuron density.
Change in Intermediate Prediction
To quantify how far the model’s representation is from the next prediction, we plot the "distance" (KL divergence) remaining for the model to reach its output distribution in Figure 6. We find that the rise in prediction neurons corresponds to a phase transition in the decrease of KL divergence. The number of prediction neurons peaks at around 85% through the model’s layers and decreases in the last few layers.
Phi-1 contains fewer prediction neurons than other Phi models and also has a lower slope in its KL divergence 6(c). GPT 6(a) and Phi 6(c) models exhibit more prediction neurons and steeper, smoother KL divergence slopes compared to Pythia 6(b). Surprisingly, Microsoft Phi models, which are known to outperform models with similar parameters, exhibit nearly 15% of prediction neurons per layer and 25% suppression neurons. This is 5-8x the density in GPT-2 and 3-7x the density in Pythia models, respectively. At around 90% through the model layers, however, the prediction neuron density decreases, while models continue approaching their final distribution, sometimes even accelerating 6(b). This suggests the action of other mechanisms, which we speculate is the final stage of inference.
@@ 5.4 Stage 4: Residual Sharpening
The subsiding of prediction neurons in the previous stage possibly suggests a new member of the ensemble, providing the final "push" to predicting the next token distribution. Our investigation reveals that the final layers of all models contain the highest density of suppression neurons, which may work to delete previously constructed features, suppress probabilities of invalid tokens, and/or calibrate the confidence in the final prediction.
Ensemble Bias
Prediction and suppression neurons both manipulate the residual stream and, as inverses of one another, can effectively perform each other’s functions. These neurons appear in different ratios and varying densities across the model. To study how these neurons sharpen the representation, we plot the logit lens entropy of the model. In certain models, such as Pythia (Figures 8(b) and 8(c)), the entropy sometimes increases in the final layers, suggesting overconfident predictions are blunted. In other words, the suppression neurons can either suppress tokens or features outside of the top-one to sharpen the distribution or suppress its confidence in the top token to flatten out the prediction distribution. This finding supports previous work, which suggests that models can shift away from the correct token to an incorrect token in the final layers [50, 61], and pruning or rank-reducing these layers can, in turn, improve performance [45, 28].
@@
The intensity of suppression neurons, as seen in Figure 6, is localized in the final few layers of the model, where the quantity of suppression neurons outstrips predictive neurons. To quantify the intensity of this change, we measure the norm of the MLP output, where a larger norm suggests a greater contribution to the residual (Figure 7). Removal of the final layer or permuting its position results in the breakage of the model (Figure 3), analogous to the breakage observed in the first layers, during which the attention norm is the greatest (Figure 5). As a result, we speculate the importance of ordering in the first and last layers due to the magnitude of change they impart.
6 Case Studies
To integrate the stages of inference hypothesis with mechanistic descriptions of models, we present two case studies. First, we identify attention heads responsible for constructing multi-token words, known as subjoiner heads [20]. These heads help capture the context of a token for appropriate prediction, thus contributing to the detokenization and feature engineering stages of models. In the second case study, we provide evidence of the ensembling of prediction and suppression neurons. Through probing experiments, we demonstrate that multiple prediction and suppression neurons working jointly significantly outperform probes trained on individual neurons and sometimes even surpass the model’s performance.
6.1 Study 1: Attention on Four-Token Words
A crucial aspect of detokenization and feature engineering is building representations that integrate the context preceding a token. For a language model to understand and predict multi-token words, it must capture both the sentence context and the tokens that comprise a single word. To study these mechanisms, we construct a dataset with two classes: each consisting of 16 tokens, where in one class, the final 4 tokens form a word. We identify specific heads in the early layers of models that contribute solely to the construction of these multi-token words. As discussed in [20], these heads are called "subjoiner" heads. As illustrated in Figure 9, layer 2 head 5 of Pythia 2.8B moves information from earlier tokens to the final token of the word. The attention heads exhibit a consistent pattern, where attention decreases as tokens approach the final word. Specifically, the final token of the word attends most strongly to the first token, a feature absent in the baseline (shown on the left in Figure 9). This suggests at least one of many mechanisms by which models integrate local context, occurring at higher density in the first half of the models.
6.2 Study 2: Predicting the suffix -ing
Neurons performing prediction ensembling must work in tandem to predict the next token - akin to voting or operating in superposition. This suggests that multiple neurons working together may form a better prediction of the next token than a single neuron. To find evidence of this mechanism, we create a balanced dataset of two classes: tokens that do or do not end with the final token of "ing", all preceded by a context of 24 tokens. We train linear probes on the activation of 32 of the most active prediction and suppression neurons, both individually and in groups. We identify these neurons as outlined by [32], and provide examples of these neurons in GPT-2 XL (Figure 10(b)). See top the 36 neurons in the Appendix 16
Probing Results
We train two types of probes on activations at the penultimate token position of the dataset. First, we train 32 individual neuron probes and measure the classification accuracy (-ing/no -ing). We compare individual probes trained with the top-k neurons of the most accurate neurons, depicted by the line in Figure 10(a). We also note the mean model accuracy when predicting a token. Probes trained on suppression neurons, shown in yellow, resulted in the highest quality individual probes and performed similarly to the model itself, depicted by the dotted red line in Figure 10(a). Top-k probes trained with prediction neurons demonstrate even better accuracy than the average model prediction accuracy. Nonetheless, an individual neuron probe performs worse than any top-k probe, suggesting a critical role for ensembling in next-token prediction.
@@ 7 Concluding Remarks
Speculations of Duality
Our findings suggest that the second half of the model is, in some sense, dual to the first half. This was briefly suggested in [15] in the context of compound words being broken down token-wise in the early layer but rebuilt in later layers. Self-repair discusses erasure and anti-erasure pairs in the first half and second half of model [44, 58], as coupled attention heads in copy suppression [43]. Zooming out, the first half of models develop complex representation and the second half has the means to clear it 6. As seen by our experiments, interference with the first layer of models is also analogous to the interference of the final layer 3. While our study only suggests this we leave it for future work to investigate this further.
Model Mystery
A mystery occurs in the final layer of the GPT model, as suggested by Figure 5. Through our finds and suggested by [54], there is attention to MLP transition in all models, except an anomalous attention spike in the final layers of GPT, indicated by the . We speculate that this is caused by tied embedding and unembedding weights ( and ) during the training of GPT-2 models. By tying these weights, GPT might be going against the “duality" discussed above. Tied weights force the “input space" and “output space" to look identical. This tying might re-involve attention units as done in the early layers of all models; however, we leave this as a future avenue to explore.
Limitations
Our study does not conclusively identify the specific causes of differences between GPT and Pythia models, such as whether redundancy stems from dropout during training, structural variations in attention and MLP mechanisms, a greater number of layers, or a combination thereof. Furthermore, the study relies on aggregation over many tokens, which may average out effects that occur to specific token classes. Despite these uncertainties, our findings suggest the existence of phases of inference across these confounding factors.
Future Work
A natural place would be to address the aforementioned model limitations and other various references throughout the text. Further investigations into the stages of inference using Sparse Autoencoders (SAEs) may form a connection to provide more evidence for or against this hypothesis.
Contributions
VL conceived and led the study, performed all the analyses, and drafted the paper. WG and MT contributed to the experimental design and analytical methodology, and provided critical revisions of the paper, with WG additionally assisting in paper writing.
Acknowledgments
Eric Michaud, Josh Engels, Dowon Baek, Isaac Liao, and the rest of the lab for helpful feedback. Thank you to MIT Supercloud for providing the compute possible for the project [57].
Appendix A Cosine Similarity Analysis of Swapped Layers
@@
Here we focus on Pythia 1.4B and GPT-2 XL, which contain a similar number of parameters (1.4B and 1.5B respectively). GPT-2 displays smoother trends compared to its Pythia counterpart while exhibiting similar overall patterns. We hypothesize that this is a result of differences in training dynamics (e.g., the use of dropout in GPT-2) and the fact that the GPT-2 model contains more layers. A larger number of layers presents greater opportunities for manipulating the output distribution and allows for more gradual changes. From an optimization perspective, this is analogous to taking smaller but more frequent gradient steps. Increasing the number of layers may also provide a means for greater redundancy in models, a key feature of GPT-style models that we discuss further below.
@@
The MLP components display two significant patterns across models. First, in the region corresponding to stage 2 of inference, we observe that the index similarity (teal line) is higher than both the adjacent similarity and the self-similarity scores. This pattern provides evidence for iterative inference, where a layer moved earlier in the computation has a projection onto the residual stream that overlaps more strongly with its previous neighbor than with its original position or its new neighbor. This overlap is more pronounced in Pythia models than in GPT-2 models, possibly because Pythia models have fewer layers to complete stage 2 of inference.
Second, in stage 3 of inference, both models demonstrate a convergence of all metrics except self-similarity toward orthogonality. The combination of high self-similarity (indicating stubbornness) and orthogonality to the replaced layer and the adjacent layers suggests a high degree of specialization in the MLPs of stage 3.
Appendix B Single Component Ablations Pythia 1.4B and GPT-2
Appendix C Top Prediction and Suppression Neurons
Appendix D Additional Empirical Details
@@
https://github.com/vdlad/Remarkable-Robustness-of-LLMs.
@@
Because of the LayerNorm operation in every layer, we can align weighs with the all-one direction in the residual stream as they do not influence the model’s calculations. Therefore, we mean-center and by subtracting the column means of :

### v2

@@ 4.3 Stage 3: Prediction Ensembling
Around the midpoint of the model, we observe a qualitative shift in behavior. Having constructed semantic features in the earlier layers, the model must begin converting these into specific next-token predictions. Evidence for this transition comes from the logit lens, where we observe a steady decline in entropy (Figure 4 right) and KL divergence (Figure 5) between intermediate and final predictions beginning around the middle layers. This suggests that the model is gradually committing to a particular output, aggregating semantic features into a more concrete distribution over tokens.
This region of the model also displays high robustness to layer interventions (Figure 2), suggesting redundancy or capacity for self-repair. One possible cause of this resilience is the presence of overlapping computational pathways [6, 51]. Rather than relying on a single deterministic path, the model seems to combine multiple signals—both across and within layers—to form its prediction. We explore this mechanism by identifying the neurons that contribute to prediction, testing their collective behavior through a case study, and analyzing their distributional effects across depth.
Ensembling
Within these overlapping pathways, we investigate specialized ensembles known as prediction neurons—units that systematically promote the likelihood of specific tokens [15, 7, 14]. These neurons work in tandem with suppression neurons (discussed in Section 4.4) to shape the model’s output.
@@
Following previous work[14], we identify these neurons by analyzing the MLP output weights and their projection into vocabulary space via the unembedding matrix . Prediction neurons exhibit a logit effect distribution with high kurtosis and positive skew; suppression neurons exhibit high kurtosis and negative skew. Across 16 models, prediction neurons begin to appear around the midpoint, increasing in density through the latter layers (Figure 5), before being overtaken by suppression neurons.
Probing for the Suffix “-ing”
We hypothesize that ensembles of prediction and suppression neurons collectively support next-token prediction. To test this, we construct a balanced classification task: given a 24-token context, does the final token end with “-ing”? We train linear probes on the activations of 32 high-variance prediction and suppression neurons, both individually and in groups. Neurons are selected using the criteria above, and examples from GPT-2 XL are shown in Figure 6. The full neuron list appears in Appendix 17.
We train two types of probes on the penultimate token’s activations: 32 individual neuron probes and top- ensemble probes ranked by individual accuracy (Figure 6). Suppression neurons yield the strongest individual probes, performing on par with the model’s predictions (dotted red line). Ensemble probes trained on prediction neurons outperform both individual neurons and the model average, suggesting an important interplay between the two neuron types.
Density Effects
The balance between prediction and suppression neurons appears to shape the model’s output. To test this, we analyze how their density relates to the KL divergence between each layer’s logit lens distribution and the final output. The sharpest decline in divergence corresponds closely with the rise in prediction neuron density, which peaks at roughly 85% of model depth.
Model comparisons further reinforce this pattern. Phi-1 has fewer prediction neurons and a shallower KL slope compared to later Phi models (Figure 5(c)). GPT and newer Phi models show steeper, smoother KL divergence drops than Pythia (Figures 5(a), 5(b)). Notably, the most performant Phi models exhibit nearly 15% prediction and 25% suppression neurons per layer—5–8 the density in GPT-2 and 3–7 that of Pythia. Interestingly, the density of prediction neurons decreases near the final 10% of layers, even as the model continues to converge on its output, sometimes accelerating(Figure 5(b)). This suggests the involvement of a distinct final-stage mechanism, which we delineate as a separate stage.
@@ 4.4 Stage 4: Residual Sharpening
(a) Probing For Neuron Ensembles in GPT2-XL
(b) MLP Output Norms
(c) Sharpening Experiment
As prediction neuron density declines in the final layers, a different mechanism appears to take over. Across all models, we observe a sharp rise in suppression neurons near the end of the network, coinciding with a general decrease in entropy (Figure 15(b)). Acting in a manner opposite to prediction neurons would suggest that suppression neurons may help refine the model’s output by removing obsolete features, down-weighting improbable tokens, and adjusting the confidence in the final prediction.
Sharpening Experiment
To further explore this hypothesis, we design an experiment where we repeat certain layers of the model. Specifically, we duplicate blocks of layers within the model—for example, repeating layers 5 through 7 results in a sequence like (…4-5-5-6-6-7-7-8-9…). For this analysis, we fix the number of repeats to 1 and the block length to 5 (see additional results in Figure 18,19). In Figure 6, we observe that repeating blocks in the latter half of the model leads to a consistent decrease in entropy relative to the baseline (horizontal line). When evaluated on downstream benchmarks, the models with repeated layers at the last 80-90% of depth also exhibit improved performance on benchmarks, probably due to clearer prediction attributed to the proposed sharpening process (Appendix 20).
@@
The intensity of suppression neurons, as seen in Figure 5, is localized in the final few layers of the model, where the quantity of suppression neurons outstrips prediction neurons. To quantify the intensity of this change to the output distribution, we measure the norm of the MLP output, where a larger norm suggests a greater contribution to the residual (Figure 6).
5 Related Work
Mechanistic Interpretability
Mechanistic interpretability often employs circuit analysis to uncover model components relevant to specific computations. In computer vision, universal mechanisms such as frequency detectors and curve-circuits have been identified [52, 53, 54], with features becoming progressively more complex through the layers of CNNs. These principles were later extended to modern transformers [55, 33], where similar circuit-based analyses revealed phenomena such as circuit reuse [56], variable-finding mechanisms [57], self-repair [6, 58], function vectors [59, 60], and long-context retrieval [61].
Iterative Inference and Depth-Dependent Computations
The iterative inference hypothesis, first explored in ResNets [62, 63], posits that each layer incrementally updates token representations. This idea has gained traction in transformers, particularly through logit lens analyses [49, 5], which visualize the model’s evolving prediction distributions layer by layer. Some studies further suggest discrete inference phases [40], with certain computations localized to specific depths—such as truth-processing [44] or multilingual translation [46]. These findings are complemented by layer permutation studies showing that performance improves when self-attention layers precede feedforward layers [64].
BERTology
Prior work on ablations and layer-wise analysis has primarily focused on BERT [65]. These studies reveal substantial redundancy: even with aggressive neuron and layer pruning, models retain most of their performance [66, 67, 68, 69, 70]. More recent investigations corroborate this, showing that a significant portion of attention heads and feedforward components can be removed with minimal accuracy loss [9, 8].
@@ 6 Concluding Remarks
Why Are Language Models Robust to Layer-Wise Interventions?
We hypothesize that the robustness of language models to layer deletion and swapping stems in part from the transformer’s residual architecture. This interpretation aligns with our findings on prediction and suppression neurons: multiple computational pathways appear to contribute to the same output, allowing the network to tolerate disruption in any single path. The residual stream promotes this “ensembling”, enabling gradient descent to construct shallow sub-networks that can operate in parallel. This architectural flexibility reduces the model’s reliance on any specific layer, explaining its resilience to local interventions and supporting observed self-repair behavior and overlapping representations.
Conclusion
This work introduces a four-stage framework for understanding inference in large language models, grounded in a diverse set of behavioral and mechanistic analyses. By examining how models respond to structural interventions—layer deletion and swapping—as well as probing attention patterns, neuron function, and residual stream dynamics, we identify a repeatable depth-wise structure to model computation. These stages—detokenization, feature engineering, prediction ensembling, and residual sharpening—emerge across architectures and scales, suggesting that transformers perform inference not as a flat pipeline but as an ordered composition of specialized computational regimes. Rather than aiming for exhaustive mechanistic dissection, we offer a unifying perspective that synthesizes and extends prior findings in and out of interpretability literature. This layered view of inference has implications for how we interpret, audit, and intervene on language models. We hope this framework serves as a foundation for a deeper investigation into the emerging capabilities of LLMs, such as reasoning.
Limitations and Future Work
While our four-stage framework captures broad, depth-dependent patterns in LLMs, several caveats remain. Stage boundaries are approximate, and multiple stages may co-occur within a single layer. The framework reflects aggregate trends, whereas individual tokens may follow distinct processing paths. Additionally, we do not isolate the factors behind model-specific differences. Future work can sharpen these boundaries, link them to optimization dynamics, and provide a more theoretical framework.
Contributions and Acknowledgements
VL conceived and led the study, performed the majority of analyses, and drafted the paper. JL conducted the sharpening experiment, WiC analysis, benchmarking, and CKA experiments. JL, WG and MT contributed to experimental design and analytical methodology, provided critical revisions, and WG and JL assisted with writing. A big thanks to Katrin Franke, Surya Ganguli, Sophia Sanborn, Eric Michaud, Josh Engels, Dowon Baek, Isaac Liao for helpful feedback.
Appendix A Appendix
A.1 Experiment Diagram
A.2 Centered Kernel Alignment (CKA)
A.3 Benchmark Tasks Performance After Layer-Wise Intervention
We evaluate the benchmark performance on HellaSwag, ARC-Easy and LAMBADA [30, 31, 32] with the intervened model. We observe a similar trend to KL divergence reported in the main paper. Generally, the intervention at the first layer and the last layer shows catastrophic deterioration of the performance but intervention on intermediate layers shows robust performance.
A.4 Cosine Similarity Analysis of Swapped Layers
@@
Here we focus on Pythia 1.4B and GPT-2 XL, which contain a similar number of parameters (1.4B and 1.5B, respectively). GPT-2 displays smoother trends compared to its Pythia counterpart while exhibiting similar overall patterns. We hypothesize that this is a result of differences in training dynamics (e.g., the use of dropout in GPT-2) and the fact that the GPT-2 model contains more layers. A larger number of layers presents greater opportunities for manipulating the output distribution and allows for more gradual changes. From an optimization perspective, this is analogous to taking smaller but more frequent gradient steps. Increasing the number of layers may also provide a means for greater redundancy in models, a key feature of GPT-style models that we discuss further below.
@@
The MLP components display two significant patterns across models. First, in the region corresponding to stage 2 of inference, we observe that the index similarity (teal line) is higher than both the adjacent similarity and the self-similarity scores. This pattern provides evidence for iterative inference, where a layer moved earlier in the computation has a projection onto the residual stream that overlaps more strongly with its previous neighbor than with its original position or its new neighbor. This overlap is more pronounced in Pythia models than in GPT-2 models, possibly because Pythia models have fewer layers to complete stage 2 of inference. Second, in stage 3 of inference, both models demonstrate a convergence of all metrics except self-similarity toward orthogonality. The combination of high self-similarity (indicating stubbornness) and orthogonality to the replaced layer and the adjacent layers suggests a high degree of specialization in the MLPs of stage 3.
A.5 Prediction and Suppression Neuron
A.6 WiC contextual word probe
A.7 Logit Lens Entropy
A.8 MLP Norms
A.9 Top Prediction and Suppression Neurons
A.10 Layer repeats experiment
A.11 Additional Empirical Details
@@
https://github.com/vdlad/Remarkable-Robustness-of-LLMs
@@
Qwen 0.5B | Qwen/Qwen2.5-0.5B |
Qwen 1.5B | Qwen/Qwen2.5-1.5B |
Qwen 3B | Qwen/Qwen2.5-3B |
Llama-3.2 1B | meta-llama/Llama-3.2-1B |
Llama-3.2 3B | meta-llama/Llama-3.2-3B |
@@
Because of the LayerNorm operation in every layer, we can align weights with the all-one direction in the residual stream as they do not influence the model’s calculations. Therefore, we mean-center and by subtracting the column means of :
Societal Impact
We do not anticipate any immediate societal impact from this research.